• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M Jones

    Log entry time 00:13:47 on May23,2002

    Entry number 4013

    keyword=Swing shift summary

    16:00 doing dummy target at 30uA runs 1779-1781
    16:45 Switch to 12C pointing at 70uA run 1782
    17:05 Switch to 4cm Lh2 runs 1783-1789
    17:45 during run 1784 fixed downstream beam v-to-f cables.
    19:30 end LH2. Make access for angle change.
    20:30 finished access.
    Setup for KIN "P" beam energy = 3.772 Q2=3.2 (left ) Q2 = 0.5 (right)
    Moved right spectrometer but magnets unchanged
    Moved left spectrometer and cycled Q2,Q3 and changed momentum.

    Beam back around 22:20 Setup intrinsic beam size
    Harp run 1710 junk
    Harp run 1711 first try 3bh = 55 3bv = 30
    harp run 1712 3bh = 54 3bv = 21 3ah = 19 3av = 38
    harp run 1713 3bh = 60 3bv = 34 = 17 = 38
    harp run 1714 = 69 = 38 = 26 = 52
    harp run 1715 = 107 = 93 = 41 = 107
    harp run 1716 129 108 55 140
    harp run 1717 = 114 = 168 = 55 = 150
    harp run 1718 = 132 = 173 = 58 = 176
    corrected position x = 0.19 mm y = -0.014mm
    Setup raster data run 1790 del(x) = 0.5 del(y) = 4.5cm